Courtney is a programmer receiving requests each week to analyze a large data base. Five jobs were on her desk Monday​ morning, and she must decide in what order to write the code. Assume that all jobs arrived today​ (day 1 and hour​ 1), in the order shown below. Courtney has assigned the number of hours required to do the coding as noted. Note that Courtney works an​ 8-hour day and today is a work day.

JOB DUE DATE (HOURS HENCE) DURATION (HOURS REQUIRED)
A 34 20
B 24 10
C 8 4
D 20 1
E 14 5

a. Sequence the jobs according to EDD?

b. What is the average competition (flow) time?

c. What is the average job lateness?

d. What is the average number of jobs in the system?
Answer #1

Answer a= Sequence as per EDD will be the ascending order of due date. So the sequence will be=C-E-D-B-A

Activity In Out Due date Lateness
C 0 4 8 0
E 4 4+5=9 14 0
D 9 9+1=10 20 0
B 10 10+10=20 24 0
A 20 20+20=40 34 6

Answer b- Averahe completion time = (4+9+10+20+40)/5=16.6

Answer c= Average job lateness = 6/5=1.2

Answer d= average number of jobs= (4+9+10+20+40)/(20+10+4+1+5)=2.08